# BiSSL: Enhancing the Alignment Between Self-Supervised Pretraining and Downstream Fine-Tuning via Bilevel Optimization

### v4

@@ C.8 Visual inspection of Latent Features
Figure 8: Features from lower-level backbones after applying BiSSL (left) or pretext pretraining (right) on the Pets [60] dataset.
@@
Figure 9: Features from lower-level backbones after applying BiSSL (left) or pretext pretraining (right) on the Aircrafts [57] dataset.
@@
Figure 10: Features from lower-level backbones after applying BiSSL (left) or pretext pretraining (right) on the Cars [56] dataset.
Appendix D Discussion of Limitations
In the interest of transparency, we highlight limitations of our current approach that may inform future work.
@@ Appendix D Discussion of Limitations
Our bilevel optimization procedure employs a conjugate gradient (CG) solver, selected for its relative simplicity and prior success in related settings. While this solver proved effective in our experiments, it may not be the most optimal in terms of computational efficiency or compatibility with the non-convex landscape inherent in deep neural networks. Future work could explore alternative solvers or reformulations of the optimization problem to further improve both accuracy and efficiency.
@@
Access to the pretext task, pretraining data (or potentially a subset thereof), and associated pretext head parameters are current requirements for applying BiSSL. This may limit applicability in some constrained environments, although future work could investigate strategies to partially decouple BiSSL from these dependencies.
@@
BiSSL also requires the backbone architectures used in the pretext and downstream stages to be identical. While this is standard in many transfer learning pipelines, it may limit flexibility in scenarios involving architectural modifications, such as the use of parameter-efficient fine-tuning methods [78]. However, this constraint could be mitigated by applying the same architectural modifications consistently across both stages of the BiSSL framework, another avenue worth exploring in future work.
@@
Finally, the benefits from BiSSL are expected to be most pronounced when there is a meaningful distributional discrepancy between the pretraining and downstream tasks. As demonstrated in Section C.7, when the downstream data distribution closely aligns with the pretext task, BiSSL may offer limited improvement. In practice, some degree of distributional shift is often present, but users should be mindful that BiSSL’s effectiveness may depend on the extent of that mismatch.

### v5

@@ C.7 Visual inspection of Latent Features
Figure 6: Features from backbones after applying self-supervised pretraining (left) and BiSSL (right) on the Pets (Parkhi et al., 2012) dataset.
@@
Figure 7: Features from backbones after applying self-supervised pretraining (left) and BiSSL (right) on the CIFAR10 (Krizhevsky, 2012) dataset.
@@
Figure 8: Features from backbones after applying self-supervised pretraining (left) and BiSSL (right) on the Aircrafts (Maji et al., 2013) dataset.
Figure 9: Features from backbones after applying self-supervised pretraining (left) and BiSSL (right) on the CUB200 (Wah et al., 2011) dataset.
Figure 10: Features from backbones after applying self-supervised pretraining (left) and BiSSL (right) on the Caltech-101 (Li et al., 2022a) dataset.
Figure 11: Features from backbones after applying self-supervised pretraining (left) and BiSSL (right) on the Cars (Yang et al., 2015) dataset.
Appendix D Discussion of Limitations and Future Work
@@
The bilevel optimization procedure of this study employs a conjugate gradient (CG) solver, selected for its relative simplicity and prior success in related settings (Rajeswaran et al., 2019; Pedregosa, 2016). While this solver proved effective in our experiments, it may not be the most optimal in terms of computational efficiency or compatibility with the non-convex landscape inherent in deep neural networks. Future work could explore alternative solvers (Huang, 2024; Hao et al., 2024; Yang et al., 2025), approximation methods for the inverse Hessian vector products (Singh and Alistarh, 2020; Frantar et al., 2021) or reformulations of the optimization problem to further improve both accuracy and efficiency.
@@
Access to the pretext task, pretraining data, and associated pretext head parameters are current requirements for applying BiSSL. This may limit applicability in some constrained environments, although future work could investigate strategies to partially or completely decouple BiSSL from these dependencies.
@@
BiSSL also requires the backbone architectures used in the pretext and downstream stages to be identical. While this is standard in many transfer learning pipelines, it may limit flexibility in scenarios involving architectural modifications, such as the use of parameter-efficient fine-tuning methods (Hu et al., 2022). However, this constraint could be mitigated by applying the same architectural modifications consistently across both stages of the BiSSL framework, another avenue worth exploring in future work.
@@
Finally, the benefits from BiSSL are expected to be most pronounced when there is a meaningful distributional discrepancy between the pretraining and downstream tasks. As demonstrated in Section C.5, when the downstream data distribution closely aligns with the pretext task, BiSSL may offer limited improvement. In practice, some degree of distributional shift is often present, but users should be mindful that the effectiveness of BiSSL may depend on the extent of that mismatch.